OpenStudy (anonymous):

Find m angle T A) 40.5° B) 76° C) 67° D) 35.4°

OpenStudy (anonymous):

picture?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi can you help?

OpenStudy (misty1212):

looks to be a law of cosines question right?

OpenStudy (misty1212):

\[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\] is how to get the cosine of the angle, then take the arc cosine to get the angle itself

OpenStudy (misty1212):

where \(a\) is the side opposite angle \(A\)

OpenStudy (misty1212):

you need a calculator, and to plug the numbers in to the right place need help with that?

OpenStudy (anonymous):

yes plz which do i put in where

OpenStudy (misty1212):

you want angle T, and the side opposite angle T has length 8.5

OpenStudy (misty1212):

so it is \[\cos(T)=\frac{13^2+11^2-8.5^2}{2\times 11\times 13}\]

OpenStudy (misty1212):

put that in the calcualtor, then take the arc cosine of the result

OpenStudy (misty1212):

or you can use wolfram

OpenStudy (misty1212):

i \(\color\magenta\heartsuit\) wolfram http://www.wolframalpha.com/input/?i=arccos%28%2811%5E2%2B13%5E2-8.5%5E2%29%2F%282*11*13%29%29

OpenStudy (misty1212):

got it?

OpenStudy (anonymous):

ok so it would be A?

OpenStudy (misty1212):

i guess

OpenStudy (misty1212):

40.5 or so

OpenStudy (anonymous):

ok thank youu

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]